OpenStudy (anonymous):

How to find log 81^1/27=x

OpenStudy (anonymous):

.0706846303 is your answer. you really need a calculator tho

OpenStudy (anonymous):

Thx

OpenStudy (triciaal):

was the question written correctly? 81 = 3^4 27 = 3^3